OpenStudy (anonymous):

write sin(5θ)-sin(θ) as a product of two trigonometric functions

OpenStudy (whpalmer4):

You can use the identity \[\sin u - \sin v = 2\cos(\frac{u+v}{2})\sin(\frac{u-v}{2})\]

OpenStudy (anonymous):

How would you write the equation using that formula?

OpenStudy (jdoe0001):

well, that is the equation in full @tigers49 http://www.sosmath.com/trig/Trig5/trig5/img9.gif

OpenStudy (jdoe0001):

just plug in your values

OpenStudy (anonymous):

oh okay thank you that's easy enough

OpenStudy (whpalmer4):

yeah, the trick is remembering the formula :-)

OpenStudy (jdoe0001):

the one whpalmer4 showed above is the so-called "sum to product" identity

OpenStudy (anonymous):

okay what about if i have something cos^24a-sin^4a as a single trigonometric function. would i use one of those formulas?

OpenStudy (anonymous):

cos^24a-sin^2 4a forgot the 2 square

OpenStudy (whpalmer4):

\[\cos^24a-\sin^24a\] Well, remember that \[\cos 2u = \cos^2u-\sin^2u\] So you can just take \(u=4a\) and write \[\cos^24a-\sin^24a=\cos 8a \]

OpenStudy (anonymous):

got it. okay so this would go also with sin6x cos2x - cos6x sin2x?

OpenStudy (whpalmer4):

You have to do some pattern matching. Here you have sin A cos B - cos A sin B. That's one of the sum/difference formulas.

OpenStudy (whpalmer4):

because each term has both a sin and cos, that's the sin sum/difference formula (for cos, the cosines are together and the sines are together). Next you have to figure out if that's the sum or the difference.

OpenStudy (whpalmer4):

With the sin sum/difference formulas, the sign on the left side matches the sign on the right side, so this is \[\sin(A-B) = \sin A \cos B - \cos A \sin B\]

OpenStudy (whpalmer4):

A = 6x, B = 2x so the final answer is?

OpenStudy (anonymous):

8x?

OpenStudy (whpalmer4):

bzzzt, wrong! but thanks for playing :-)

OpenStudy (whpalmer4):

sin(A-B), A = 6x, B = 2x

OpenStudy (whpalmer4):

A-B =

OpenStudy (anonymous):

sin4x

OpenStudy (whpalmer4):

now we're talking!

OpenStudy (whpalmer4):

do you have a good list of trig identities?

OpenStudy (whpalmer4):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html is the one I use

OpenStudy (anonymous):

No i don't.

OpenStudy (whpalmer4):

make yourself a set of flashcards on 3x5 cards and drill on them for 5 minutes a day. it'll be time well spent in the long run...

OpenStudy (anonymous):

Thank you I'm sure I'll be using this more real soon again.